OpenStudy (anonymous):

08.04a Parametric Equations: Graph each pair of parametric equations. x = 2t y = t + 5; -2 ≤ t ≤ 3 Please help! I am not sure how to solve this at all, I have tried everything! Medal will be awarded to who helps me!

OpenStudy (anonymous):

@amistre64 can you help me?

OpenStudy (anonymous):

@TuringTest

OpenStudy (amistre64):

well, there are a few ways to approach this, one is to solve t for x, and sub in the y part; this creates a more usual y = f(x) format

OpenStudy (amistre64):

the other is to table it x y t: 2t : t+5 -2: -4 : 3 -1: -2 : 4 0: 0 : 5 1: 2 : 6 2: 4 : 7 3: 6 : 8

OpenStudy (amistre64):

x = 2t ; t = x/2 2(-2) <= x <= 2(3) -4 <= x <= 6 y = t + 5 ; y = x/2 + 5 is a simple enough line to graph

OpenStudy (anonymous):

Thank you so much!!